An ideal transformer has a primary winding with 100 turns and a secondary winding with 20 turns. If the primary winding has a voltage drop of 10 V AC, what is the EMF in the secondary coil? A. 100 V AC B. 10 V AC C. 2 V AC D. 20 V AC
Answer #1

Transformer formula

\frac{V_{p}}{V_{s}}=\frac{N_{p}}{N_{s}}

where V is voltage or EMF and N is number of turns.

which gives

{V_{s}}=V_{p}\frac{N_{s}}{N_{p}}=\frac{10\times 20}{100}=2 V AC

Option c.

I will be glad to see your comment if you have any query and thumb up if you are satisfied. Thanks...